solve for W in the equation XW + Y = B / 4
show restrictions:

Answer by nerdybill(7384)

Simply isolate W to one side of the equation:
XW + Y = B/4
XW = (B/4) - Y
W = (B/4X) - (Y/X)
.
The only restriction is:
Except when X=0
.
Notice that if X was equal 0, the numerator would be zero and a zero in the denominator would be "undefined".